find two numbers who sum is -66 and whose difference is -116
Answer by algebrahouse.com(1659)

"find two numbers who sum is -66 and whose difference is -116"

x = one number
y = other number

x + y = -66 {sum is -66}
x - y = -116 {difference is -116}
------------
2x = -182 {added the two equations together}
x = -91 {divided both sides by 2}

x + y = -66 {top equation}
-91 + y = -66 {substituted -91, in for x, into top equation}
y = 25 {added 91 to both sides}

Two numbers are -91 and 25

*To check*
-91 + 25 = -66 {sum is -66}
-91 - 25 = -116 {difference is -116}
www.algebrahouse.com